OpenStudy (anonymous):

insert five(5) geometric means between : 1) 4 & 49 2) square root of 3 & 9 3) 1 over two & 32 4) 2 & 54 i need this to my homework .. please Help me .. :)) thanks .

OpenStudy (anonymous):

I am not sure what is meant by "insert five", but the geometric mean of two numbers \(a\) and \(b\) is given by \(\sqrt{ab}\).

OpenStudy (anonymous):

im sorry i dont know too .. my adviser just give the problem to us ..

OpenStudy (anonymous):

Well, two numbers cannot have more than one geometric mean. I'd just evaluate the geometric means of the given pairs of numbers.

OpenStudy (anonymous):

ok .

OpenStudy (anonymous):

ok .

OpenStudy (anonymous):

please read this . where are u ?

OpenStudy (anonymous):

I can't help you unless you ask a question.

OpenStudy (anonymous):

ok . i mean its not geometric .. its arithmetic .

OpenStudy (anonymous):

The arithmetic mean is \[\frac{a+b}{2}\]whereas the geometric mean is \[\sqrt{ab}.\]Either way, it's just evaluating the formula.

OpenStudy (anonymous):

OMG how can answer it i dont know it ......... thanks to your help ..

OpenStudy (anonymous):

Let me give you an example. Take the numbers 2 and 3. Their arithmetic mean is\[\frac{2+3}{2} = \frac{5}{2}.\]Their geometric mean is \[\sqrt{2 \cdot 3} = \sqrt{6}.\]

OpenStudy (anonymous):

my adviser teach us arithmetic like this ............ the difference is equals to 3 : 3, 6 , 9 , ... and so on .. what do u think ?

OpenStudy (anonymous):

That is an arithmetic series, which is distinct from an arithmetic mean. You wrote geometric mean in your question.

OpenStudy (anonymous):

ohh .. yes i remember its arithmetic series , thanks for remind me :))

OpenStudy (anonymous):

can u answer it .. ??

OpenStudy (anonymous):

I still do not understand what you are asking. The question you wrote doesn't make sense.

OpenStudy (anonymous):

ok . i dont have my chance ......

OpenStudy (anonymous):

i dont know it .. i PROMISE !!!!

OpenStudy (whpalmer4):

@yakeyglee Here's what I think they mean: for example, insert 3 means between 2 and 32 would mean the result is 4, 8, 16

OpenStudy (anonymous):

yes .. your right @whpalmer4 .

OpenStudy (anonymous):

thats what i mean so .

OpenStudy (anonymous):

hey !!

OpenStudy (whpalmer4):

Okay, so if you have two numbers \(x,y\) and you want to insert \(n\) geometric means, you find the common ratio by taking \[\frac yx = a^{n+1}\] and solving for \(a\). For my example of 3 means between 2 and 32, \[\frac{32}{2} = a^{3+1}\]\[16=a^4\]\[a=2\]So our numbers are 2, 4, 8, 16, 32

OpenStudy (whpalmer4):

For the first problem, \[\frac{49}{4} = a^{5+1}\]\[a=\sqrt[6]{\frac{49}4}\approx 1.51829\]So the values of the sequence would be 4,6.07318, 9.22087, 14, 21.2561, 32.2731, 49